# FasterCache: Training-Free Video Diffusion Model Acceleration with High Quality

### v1

@@ Abstract
footnotetext: †Corresponding authors. $\ddagger$ Project leader. *The work was done during an internship at Shanghai AILab.
@@ 1 Introduction
Diffusion transformers (DiT) (Peebles & Xie, 2023) have achieved notable success in image (Chen et al., 2023; 2024b; Esser et al., 2024) and video generation (Ma et al., 2024a; Zheng et al., 2024; Lab & etc., 2024), attracting significant attention for their potential. Although iterative denoising, classifier-free guidance (CFG) (Ho & Salimans, 2022), and transformer attention mechanisms have significantly improved the generative capabilities of diffusion models, they also lead to substantial computational costs and increased memory requirements for inference, especially for video generation which typically takes 2-5 minutes to synthesize a 6-second 480P video, limiting their practical use. This calls for the development of new techniques that require less computational cost for diffusion models (Salimans & Ho, 2022; Ma et al., 2024b; Chen et al., 2024c; Zhao et al., 2024c).
Among the recently proposed solutions, cache-based acceleration has emerged as one of the most widely adopted approaches. This approach speeds up the sampling process by reusing intermediate features across timesteps, thereby reducing redundant computations and significantly improving computational efficiency. Besides, it requires no additional training costs for inference acceleration and offers straightforward generalization to other video diffusion models. Examples include the cache-based methods for U-Net based diffusion models (Ma et al., 2024b; Li et al., 2023b), residual caching in $\Delta$-DiT (Chen et al., 2024c) for the transformer based diffusion models, and hierarchical attention caching of PAB (Zhao et al., 2024c) for video generation. Despite their proven effectiveness, there exist two critical concerns: 1) Whether directly reusing intermediate features aligns with the iterative denoising mechanism, considering the inherent feature variations between timesteps. 2) Current cache-based methods focus primarily on the attention features within the transformer networks, with limited exploration of accelerating different parts of the pipeline. In this work, we aim to address these two concerns.
Figure 2: Vanilla cache-based acceleration method.
@@
We evaluate our FasterCache on various video diffusion models, including Open-Sora 1.2 (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), Latte Ma et al. (2024a), CogVideoX (Yang et al., 2024), and Vchitect-2.0 (Vchitect, 2024). Experimental results demonstrate that FasterCache can significantly accelerate inference while preserving high-quality video generation across all tested models. Specifically, on Vchitect-2.0, FasterCache achieves 1.67$\times$ speedup, with performance comparable to the baseline (VBench: baseline 80.80% $\rightarrow$ FasterCache 80.84%). Furthermore, our method outperforms existing approaches in both inference speed and video generation quality, highlighting its effectiveness and efficiency in real-world applications.
@@ 2.1 Preliminary
Video diffusion models recently employ diffusion transformers as the backbone for noise prediction. This work explores video synthesis acceleration based on Open-Sora 1.2 (Zheng et al., 2024). This model is composed of 56 stacked transformer layers, with alternating spatial and temporal layers. Each layer contains not only a spatial or temporal attention module but also a cross-attention and a feed-forward network. Latte (Ma et al., 2024a) and Open-Sora-Plan (Lab & etc., 2024) also adopt a similar architecture as their noise prediction networks.
@@ 2.2 Rethinking Attention Feature Reuse
To investigate the underlying causes of this phenomenon, we subsequently visualize the attention features between adjacent timesteps and analyze their differences. The results indicate that while the attention features between adjacent timesteps are highly similar, there exist noticeable differences between them. These subtle variations between timesteps are essential for preserving fine details in video generation. Therefore, directly reusing features without accounting for these differences leads to the loss of important visual information, resulting in smoother but less detailed outputs. This highlights the need for a more refined approach to feature reuse, i.e., one that can retain computational efficiency while preserving key inter-step variations.
@@ 2.4 FasterCache for Video Diffusion model
where $w(t)$ is a weighting function that modulates the contribution of the feature difference to account for variation between adjacent timesteps, ensuring both efficiency and the preservation of fine details in the generated videos. In our experiments, $w(t)$ gradually increases as the sampling process progresses, allowing the model to place greater emphasis on the feature differences at later stages of generation. Consequently, our approach significantly accelerates inference while preserving the visual quality of the synthesized videos.
CFG-Cache As analyzed in Section 2.3, the conditional and unconditional outputs at the same timestep exhibit high similarity in CFG, indicating significant information redundancy. A naive approach to take advantage of this would be to directly reuse the conditional features for the corresponding unconditional outputs at the same timestep. However, this often leads to a noticeable degradation in detail generation. As illustrated in Fig 7 (a), this approach results in poor generation of intricate details, such as the texture of the spacesuit which shows a lack of details and clarity. Since both the conditional and unconditional outputs in CFG represent predicted noise, and drawing inspiration from the Dynamic Feature Reuse Strategy and FreeU (Si et al., 2024), we analyze the differences between these two outputs in the frequency domain. As shown in Fig 7 (b), we observe that, in the early and mid-stages of the sampling process, the conditional and unconditional outputs at the same timestep exhibit a significant bias in the low-frequency components, which progressively shifts to the high-frequency components in the later steps. This observation suggests that despite their overall similarity, key differences in frequency components must be addressed to avoid the degradation of fine details.
Figure 8: Overview of the CFG-Cache.
@@
$\Delta_{LF}=\mathcal{FFT}(\bm{\epsilon}_{\theta}(\bm{x}_{t},t,\emptyset))_{low}-\mathcal{FFT}(\bm{\epsilon}_{\theta}(\bm{x}_{t},t,\bm{c}))_{low}$ (6)
@@
$\mathcal{F}_{low}$ $=\quad\Delta_{LF}*w_{1}+\mathcal{FFT}(\bm{\epsilon}_{\theta}(\bm{x}_{t-i},t-i,\bm{c}))_{low}$ (9)
$\mathcal{F}_{high}$ $=\Delta_{HF}*w_{2}+\mathcal{FFT}(\bm{\epsilon}_{\theta}(\bm{x}_{t-i},t-i,\bm{c}))_{high}$ (10)
@@ 3.1 Experimental Settings
Base models and compared methods To demonstrate the effectiveness of our method, we apply our acceleration technique to various video diffusion models, including Open-Sora 1.2 (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), Latte (Ma et al., 2024a), CogVideoX (Yang et al., 2024), and Vchitect-2.0 (Vchitect, 2024). We compare our base models with recent efficient video synthesis techniques, including PAB (Zhao et al., 2024c) and $\Delta$-DiT (Chen et al., 2024c), to highlight the benefits of our approach. Notably, $\Delta$-DiT was originally designed as an acceleration method for image synthesis. Here we have adapted it for video synthesis to facilitate comparison. Please refer to the Appendix for more details of the base models and compared methods.
Evaluation metrics and datasets To assess the performance of video synthesis acceleration methods, we focus primarily on two aspects, namely inference efficiency and visual quality. To evaluate inference efficiency, we employ Multiply-Accumulate Operations (MACs) and inference latency as metrics. We utilize VBench (Huang et al., 2024), LPIPS (Zhang et al., 2018), PSNR, and SSIM for visual quality evaluation. VBench is a comprehensive benchmark suite for video generative models. It is well-aligned with human perceptions and capable of providing valuable insights from multiple perspectives. LPIPS, PSNR, and SSIM measure the similarity between videos generated by the accelerated sampling method and those from the original model. PSNR quantifies pixel-level fidelity between outputs, LPIPS measures perceptual consistency, and SSIM assesses structural similarity. In general, higher similarity scores indicate better fidelity and visual quality.
Implementation details All experiments conduct full attention inference for spatial and temporal attention modules every 2 timesteps to facilitate dynamic feature reuse. The weight $w(t)$ increases linearly from 0 to 1 starting from the beginning of dynamic feature reuse until the end of sampling. For CFG output reuse, full inference is conducted every five timesteps, starting from $1/3$ of the total sampling steps (e.g., for Open-Sora 1.2, which has 30 total sampling steps, this begins at step 10). The hyperparameters $\alpha_{1}$ and $\alpha_{2}$ are set to a default value of 0.2, which performs well for most models. All experiments are carried out on NVIDIA A100 80GB GPUs using PyTorch, with FlashAttention (Dao et al., 2022) enabled by default.
@@ 3.2 Main Results
Figure 9: Comparison of visual result quality among different methods.
@@ 3.3 Ablation Study
Variants VBench LPIPS SSIM PSNR
@@ 4.1 Diffusion Models for Video Synthesis
Diffusion models have demonstrated potential in high-quality image synthesis (Ho et al., 2020; Rombach et al., 2022; Chen et al., 2023; 2024b), attracting significant attention. Subsequent work has adapted these models for video synthesis to generate high-fidelity videos (Ho et al., 2022). Motivated by advancements in image synthesis, early studies typically employed the diffusion UNet architecture (Blattmann et al., 2023; Wang et al., 2023; Zhang et al., 2023; Wu et al., 2023). As the scalability of diffusion transformer (Peebles & Xie, 2023) was validated in image synthesis, an increasing number of works have adopted the diffusion transformer as the noise estimation network (Ma et al., 2024a; Zheng et al., 2024; Lab & etc., 2024; Yang et al., 2024).
@@ 4.2 Efficiency Improvements in Diffusion Models
Despite the impressive performance of diffusion models in image and video synthesis, their substantial inference cost limits their practicality. Prior research on improving the efficiency of diffusion models has primarily focused on two perspectives, namely reducing the number of sampling steps and lowering the inference cost per sampling step. Regarding the reduction of sampling steps, most approaches achieve high-quality samples with fewer sampling steps by employing efficient SDE or ODE solvers (Song et al., 2020; Lu et al., 2022a; b). Other methods reduce the number of sampling steps by progressively distilling the model (Salimans & Ho, 2022; Sauer et al., 2023; Lin & Yang, 2024; Li et al., 2024b) or employing consistency models (Luo et al., 2023; Song et al., 2023).
@@ A.1 Further Details of base models
In this work, we applied our FasterCache to various video synthesis models, including Open-Sora 1.2 (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), Latte (Ma et al., 2024a), CogVideoX (Yang et al., 2024), and Vchitect 2.0 Vchitect (2024). Open-Sora 1.2 (Zheng et al., 2024) integrates 2D-VAE and 3D-VAE to enhance video compression and employs ST-DiT blocks for the diffusion process. Open-Sora-Plan (Lab & etc., 2024) adopts CausalVideoVAE to compress visual representations better and 3D full attention architecture to capture joint spatial and temporal features. Latte Ma et al. (2024a) extracts spatio-temporal tokens from input videos and then adopts a series of transformer blocks to model video distribution in the latent space. CogVideoX (Yang et al., 2024) employs a 3D VAE to compress videos along spatial and temporal dimensions and an expert transformer with the expert adaptive LayerNorm to facilitate the fusion between the two modalities.
@@ A.2 Further Details of compared methods
A.3 Additional Qualitative Experiments

### v2

@@ Abstract
[…]
@@ 1 Introduction
Diffusion transformers (DiT) (Peebles & Xie, 2023) have achieved notable success in image (Chen et al., 2023; 2024b; Esser et al., 2024) and video generation (Ma et al., 2024a; Zheng et al., 2024; PKU-Yuan Lab and Tuzhan AI etc., 2024), attracting significant attention for their potential. Although iterative denoising, classifier-free guidance (CFG) (Ho & Salimans, 2022), and transformer attention mechanisms have significantly improved the generative capabilities of diffusion models, they also lead to substantial computational costs and increased memory requirements for inference, especially for video generation which typically takes 2-5 minutes to synthesize a 6-second 480P video, limiting their practical use. This calls for the development of new techniques that require less computational cost for diffusion models (Salimans & Ho, 2022; Ma et al., 2024b; Chen et al., 2024c; Zhao et al., 2024c).
Among the recently proposed solutions, cache-based acceleration has emerged as one of the most widely adopted approaches. This approach speeds up the sampling process by reusing intermediate features across timesteps, thereby reducing redundant computations and significantly improving computational efficiency. Besides, it requires no additional training costs for inference acceleration and offers straightforward generalization to other video diffusion models. Examples include cache-based methods for U-Net based diffusion models (Ma et al., 2024b; Li et al., 2023b), residual caching in $\Delta$-DiT (Chen et al., 2024c) for transformer based diffusion models, and hierarchical attention caching of PAB (Zhao et al., 2024c) for video generation. Despite their proven effectiveness, there exist two critical concerns: 1) Whether directly reusing intermediate features aligns with the iterative denoising mechanism, considering the inherent feature variations between timesteps. 2) Current cache-based methods focus primarily on the attention features within the transformer networks, with limited exploration of accelerating different parts of the pipeline. In this work, we aim to address these two concerns.
Figure 2: Vanilla cache-based acceleration method typically reuses features cached from previous timesteps directly for the current timestep.
@@
We evaluate our FasterCache on various video diffusion models, including Open-Sora 1.2 (Zheng et al., 2024), Open-Sora-Plan (PKU-Yuan Lab and Tuzhan AI etc., 2024), Latte (Ma et al., 2024a), CogVideoX (Yang et al., 2024), and Vchitect-2.0 (Fan et al., 2025). As shown in Fig 1, experimental results demonstrate that FasterCache can significantly accelerate inference while preserving high-quality video generation across all tested models. Specifically, on Vchitect-2.0, FasterCache achieves 1.67$\times$ speedup, with performance comparable to the baseline (VBench: baseline 80.80% $\rightarrow$ FasterCache 80.84%). Furthermore, our method outperforms existing approaches in both inference speed and video generation quality, highlighting its effectiveness and efficiency in real-world applications.
@@ 2.1 Preliminary
Video diffusion models recently employ diffusion transformers as the backbone for noise prediction. This work explores video synthesis acceleration based on Open-Sora 1.2 (Zheng et al., 2024). This model is composed of 56 stacked transformer layers, with alternating spatial and temporal layers. Each layer contains not only a spatial or temporal attention module but also a cross-attention and a feed-forward network. Latte (Ma et al., 2024a) and Open-Sora-Plan (PKU-Yuan Lab and Tuzhan AI etc., 2024) also adopt a similar architecture as their noise prediction networks.
@@ 2.2 Rethinking Attention Feature Reuse
To investigate the underlying causes of this phenomenon, we subsequently visualize the attention features between adjacent timesteps and analyze their differences. The results indicate that while the attention features between adjacent timesteps are highly similar, there exist noticeable differences between them. These subtle variations between timesteps are essential for preserving fine details in video generation. Therefore, directly reusing features without accounting for these differences leads to the loss of important visual information, resulting in smoother and less detailed outputs. This highlights the need for a more refined approach to feature reuse, i.e., one that can retain computational efficiency while preserving key inter-step variations.
@@ 2.4 FasterCache for Video Diffusion model
where $w(t)$ is a weighting function that modulates the contribution of the feature difference to account for variation between adjacent timesteps, ensuring both efficiency and the preservation of fine details in the generated videos. In our experiments, $w(t)$ gradually increases as the sampling process progresses, allowing the model to place greater emphasis on the feature differences at later stages of generation. Further discussions on the design of feature bias term and the selection of $w(t)$ in Eq. (5) can be found in Appendix A.3.1. Consequently, our approach significantly accelerates inference while preserving the visual quality of the synthesized videos.
CFG-Cache As analyzed in Section 2.3, the conditional and unconditional outputs at the same timestep exhibit high similarity in CFG, indicating significant information redundancy. A naive approach to take advantage of this would be to directly reuse the conditional features for the corresponding unconditional outputs at the same timestep. However, this often leads to a noticeable degradation in detail generation. As illustrated in Fig 7 (a), this approach results in poor generation of intricate details, such as the texture of the spacesuit which shows a lack of details and clarity. Since both the conditional and unconditional outputs in CFG represent predicted noise, and drawing inspiration from the Dynamic Feature Reuse Strategy and FreeU (Si et al., 2024), we analyze the differences between these two outputs in the frequency domain. In Fig 7 (b), we observe that, from the activation of CFG-Cache until the end of the sampling, the difference between the conditional and unconditional outputs gradually shifts from being dominated by low-frequency components to being dominated by high-frequency components. This indicates that the effects of CFG in the sampling process is primarily to influence perceptual features like layout and shape during the early and mid-stages, while contributing to detail synthesis in the later stages. A similar phenomenon can also be observed in  Hsiao et al. (2024). This observation suggests that despite their overall similarity, key differences in frequency components must be addressed to avoid the degradation of fine details. More discussion and visualization can be found in Appendix A.3.2.
Figure 8: Overview of the CFG-Cache. CFG-Cache accelerates the computation of the unconditional output (in the dashed orange box) by caching the high- and low-frequency biases between the conditional and unconditional outputs, and dynamically enhancing them during reuse.
@@
$\Delta_{LF}$ $=\mathcal{FFT}(\bm{\epsilon}_{\theta}(\bm{x}_{t},t,\emptyset))_{low}-\mathcal{FFT}(\bm{\epsilon}_{\theta}(\bm{x}_{t},t,\bm{c}))_{low},$ (6)
@@
$\mathcal{F}_{low}$ $=\quad\Delta_{LF}*w_{1}+\mathcal{FFT}(\bm{\epsilon}_{\theta}(\bm{x}_{t-i},t-i,\bm{c}))_{low},$ (9)
$\mathcal{F}_{high}$ $=\Delta_{HF}*w_{2}+\mathcal{FFT}(\bm{\epsilon}_{\theta}(\bm{x}_{t-i},t-i,\bm{c}))_{high}.$ (10)
@@ 3.1 Experimental Settings
Base models and compared methods To demonstrate the effectiveness of our method, we apply our acceleration technique to different video synthesis diffusion models, including the Open-Sora 1.2 (Zheng et al., 2024), Open-Sora-Plan (PKU-Yuan Lab and Tuzhan AI etc., 2024), Latte (Ma et al., 2024a), CogVideoX (Yang et al., 2024), and Vchitect-2.0 (Fan et al., 2025). We compare our base models with recent efficient video synthesis techniques, including PAB (Zhao et al., 2024c) and $\Delta$-DiT (Chen et al., 2024c), to highlight the benefits of our approach. Notably, $\Delta$-DiT was originally designed as an acceleration method for image synthesis. Here we have adapted it for video synthesis to facilitate comparison. Please refer to the Appendix for more details of the base models and compared methods.
Evaluation metrics and datasets To assess the performance of video synthesis acceleration methods, we focus primarily on two aspects, namely inference efficiency and visual quality. To evaluate inference efficiency, we employ Multiply-Accumulate Operations (MACs) and inference latency as metrics. We utilize VBench (Huang et al., 2024), LPIPS (Zhang et al., 2018), PSNR, and SSIM for visual quality evaluation. VBench is a comprehensive benchmark suit for video generative models. It is well-aligned with human perceptions and capable of providing valuable insights from multiple perspectives. LPIPS, PSNR, and SSIM measure the similarity between videos generated by the accelerated sampling method and those from the original model. PSNR quantifies pixel-level fidelity between outputs, LPIPS measures perceptual consistency, and SSIM assesses structural similarity. In general, higher similarity scores indicate better fidelity and visual quality.
Implementation details All experiments conduct full attention inference for spatial and temporal attention modules every 2 timesteps to facilitate dynamic feature reuse. The weight $w(t)$ increases linearly from 0 to 1 starting from the beginning of dynamic feature reuse until the end of sampling. For CFG output reuse, full inference is conducted every 5 timesteps, starting from $1/3$ of the total sampling steps (e.g., for Open-Sora 1.2, which has 30 total sampling steps, this begins at step 10). The hyperparameters $\alpha_{1}$ and $\alpha_{2}$ are set to a default value of 0.2, which performs well for most models. For more details on the selection of hyperparameters, please refer to Appendix A.5. All experiments are carried out on NVIDIA A100 80GB GPUs using PyTorch, with FlashAttention (Dao et al., 2022) enabled by default.
@@ 3.2 Main Results
Figure 9: Visual quality comparison of different methods. Differences are highlighted in red boxes.
@@ 3.3 Ablation Study
Variants VBench LPIPS PSNR SSIM
@@ 4.1 Diffusion Models for Video Synthesis
Diffusion models have demonstrated potential in high-quality image synthesis (Ho et al., 2020; Rombach et al., 2022; Chen et al., 2023; 2024b), attracting significant attention. Subsequent works have adapted these models for video synthesis to generate high-fidelity videos (Ho et al., 2022). Motivated by advancements in image synthesis, early studies typically employed the diffusion UNet architecture (Blattmann et al., 2023; Wang et al., 2023; Zhang et al., 2023; Wu et al., 2023; Zhang et al., 2024c). As the scalability of diffusion transformer (Peebles & Xie, 2023) was validated in image synthesis, an increasing number of works have adopted the diffusion transformer as the noise estimation network (Ma et al., 2024a; Zheng et al., 2024; PKU-Yuan Lab and Tuzhan AI etc., 2024; Yang et al., 2024).
@@ 4.2 Efficiency Improvements in Diffusion Models
Despite the impressive performance of diffusion models in image and video synthesis, their substantial inference cost limits their practicality. Prior research on improving the efficiency of diffusion models has primarily focused on two perspectives, namely reducing the number of sampling steps and lowering the inference cost per sampling step. Regarding the reduction of sampling steps, most approaches achieve high-quality samples with fewer steps by employing efficient SDE or ODE solvers (Song et al., 2020; Lu et al., 2022a; b). Other methods reduce sampling steps by progressively distilling the model (Salimans & Ho, 2022; Meng et al., 2023; Sauer et al., 2023; Lin & Yang, 2024; Li et al., 2024b) or employing consistency models (Luo et al., 2023; Song et al., 2023).
@@ 5 CONCLUSION AND DISCUSSION
6 Acknowledgements
This study is supported by the National Key R&D Program of China No.2022ZD0160102, and by the video generation project (Intern-Vchitect) of Shanghai Artificial Intelligence Laboratory. This study is also supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221-0012, MOE-T2EP20223-0002), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ A.1 Further Details of base models
In this work, we applied our FasterCache to various video synthesis models, including Open-Sora 1.2 (Zheng et al., 2024), Open-Sora-Plan (PKU-Yuan Lab and Tuzhan AI etc., 2024), Latte (Ma et al., 2024a), CogVideoX (Yang et al., 2024), and Vchitect 2.0 (Fan et al., 2025). Open-Sora 1.2 integrates 2D-VAE and 3D-VAE to enhance video compression and employs ST-DiT blocks for the diffusion process. Open-Sora-Plan adopts CausalVideoVAE to compress visual representations better and 3D full attention architecture to capture joint spatial and temporal features. Latte extracts spatio-temporal tokens from input videos and then adopts a series of transformer blocks to model video distribution in the latent space. CogVideoX employs a 3D VAE to compress videos along spatial and temporal dimensions and an expert transformer with the expert adaptive LayerNorm to facilitate the fusion between the two modalities.
@@ A.2 Further Details of compared methods
A.3 More Discussion
A.3.1 More Discussion on Dynamic Feature Reuse
Effectiveness of Dynamic Feature Reuse Assume that the output features of a particular layer in the diffusion model are a function of the timestep $t$, denoted as $F(t)$. The motivation behind Vanilla Feature Reuse lies in the observation that features at adjacent timesteps are highly similar. Vanilla Feature Reuse avoids the computation at the current timestep by directly reusing the features from the previous timestep, i.e. $F(t)=F(t+\Delta t)$. Although $F(t)$ and $F(t+\Delta t)$ are very close with a minimal error $E=F(t)-F(t+\Delta t)$, the difference is not zero. To estimate this error, we assume that $F(t)$ is a smooth and differentiable function with respect to $t$, allowing us to perform a Taylor expansion, yielding:
$F(t+\Delta t)=F(t)+\frac{dF(t)}{dt}\Delta t+\frac{d^{2}F(t)}{dt^{2}}\frac{\Delta t^{2}}{2}+O(\Delta t^{3}),$ (12)
$F(t+3\Delta t)=F(t)+3\frac{dF(t)}{dt}\Delta t+3\frac{d^{2}F(t)}{dt^{2}}\frac{\Delta t^{2}}{2}+O(\Delta t^{3}).$ (13)
By subtracting these expansions, we derive:
$F(t+\Delta t)-F(t+3\Delta t)=(\frac{dF(t)}{dt}\Delta t)\times(-2)+O(\Delta t^{2}),$ (14)
Based on the statistics of approximately 200 video samples, we plotted the magnitudes of the first-order and second-order terms of $F(t)$. When $\Delta t$ (e.g., $\Delta t=1$) is sufficiently small, the norm of second-order term is smaller than that of the first-order term, as shown in Fig. 13 (c). Furthermore, we tested three different estimations for $F(t)$, denoted as $\hat{F}(t)$: (a) $\hat{F}(t)=F(t+1)$, (b) $\hat{F}(t)=F(t+1)-\frac{dF(t)}{dt}$, and (c) $\hat{F}(t)=F(t+1)-\frac{dF(t)}{dt}-\frac{d^{2}F(t)}{2dt^{2}}$. Subsequently, we calculated the L1 distance between each $\hat{F}(t)$ and $F(t)$. As shown in the Fig. 13 (d), incorporating the second-order term yields only a marginal reduction in the L1 distance compared to the first-order term. Therefore, the second-order terms contribute only marginally to the improvement in visual quality (VBench: 78.77% $\rightarrow$ 78.80%). However, the computation of second-order terms incurs significant costs in memory and latency. Considering both simplicity and efficiency, we use only the first-order term for error estimation in Dynamic Feature Reuse. Based on these analyses and statistical results, we define the error term as:
$E=F(t)-F(t+\Delta t)\approx-\frac{dF(t)}{dt}\Delta t=(F(t+\Delta t)-F(t+3\Delta t))*w.$ (15)
The scale factor $w$ is introduced to scale the bias term to approximate the error $E$. In Eq. (5), $E=F_{t-1}-F_{cache}^{t}\approx(F_{cache}^{t}-F_{cache}^{t+2})*w(t)$. By introducing this feature bias term, the information loss could be reduced, thereby improving the quality of the synthesis videos while maintaining computational efficiency.
Design choices for $w(t)$ in Dynamic Feature Reuse As shown in Fig. 13 (a), we tried different design choices for Dynamic Feature Reuse (DFR) and found that the linear increasing strategy is a simple and effective manner for dynamically capturing missing features. Different design choices for DFR: (1) Constant weights $w(t)$. A constant weight of $w(t)=0.5$ is applied to the feature biases at each accelerated timesteps. (2) Learnable weights $w(t)$. We introduced a set of learnable parameters $w(t)$, which are optimized by minimizing the MSE loss between the features output by DFR during accelerated sampling and those generated in the original unaccelerated sampling process, resulting in the learned $w(t)$. (3) Linearly increasing $w(t)$ (Our DFR). Starting from the application of DFR to the end of sampling proces, the weight function $w(t)$, used for weighting feature biases, linearly increases from 0 to 1.
The trend of the optimized $w(t)$ is shown in Fig 13 (a), the result indicates that $w(t)$ obtained through optimization gradually increases as sampling progresses. This trend is primarily attributed to the increasing stability of feature biases in Eq. 5 with respect to the sampling timesteps and the growing reliance on bias features for synthesizing high-quality details in the later stages of sampling. The performance of different strategies is shown in Table 5. All results incorporating feature biases outperform those without them. The linearly increasing $w(t)$ achieves comparable performance to optimized learnable $w(t)$, both outperforming constant $w(t)$. Given the simplicity of linear interpolation, we ultimately adopt linearly interpolated $w(t)$ to weight the feature biases.
Figure 13: Design choices for Dynamic Feature Reuse and comparison between Dynamic Feature Reuse (DFR) and Vanilla Feature Ruse (VFR).
Table 5: Performance of different Dynamic FR strategies.
Variants LPIPS PSNR SSIM
Vanilla FR 0.0657 28.20 0.8785
Dynamic FR (Constant $w(t)=0.5$) 0.0615 28.33 0.8889
Dynamic FR(Learned $w(t)$) 0.0596 28.45 0.8941
Dynamic FR(Linear $w(t)$) 0.0590 28.41 0.8938
Comparison between Dynamic FR and Vanilla FR Fig. 13 (b) presents the generated results of Vanilla Feature Reuse (FR) and Dynamic FR and the differences between the features produced by Vanilla FR and Dynamic FR compared to the original features. It is evident that, due to the introduction of feature biases, the feature differences between Dynamic FR and the original features are less significant. In contrast, the features produced by the model accelerated with Vanilla FR exhibit detail loss compared to the original features, leading to noticeable detail degradation in the synthesized images (as highlighted by the red box).
A.3.2 Further discussion on CFG-Cache
Effectiveness of CFG-Cache The reliability of CFG-Cache stems from three key factors: (a) After the early stage $t_{early}$, the similarity between conditional output $cond(t)$ and unconditional output $uncond(t)$ at the same timestep $t$:
$uncond(t)=cond(t)+\Delta,when\ t>=t_{early}.$ (16)
(b) The predictability of biases between conditional and unconditional output from previous timesteps, expressed as:
$\Delta=uncond(t+\Delta t)-cond(t+\Delta t)=uncond(t)-cond(t)+\epsilon.$ (17)
In practice, we find that when $\Delta t$ is sufficiently small, the $\epsilon$ can be considered negligible. Then:
$uncond(t)\approx cond(t)+(uncond(t+\Delta t)-cond(t+\Delta t))$ (18)
(c) The dynamic variations of the frequency-domain distribution of feature biases, as illustrated in Fig. 7(b) and Fig.14.
Visualization of CFG biases From the onset of CFG-Cache to the end of sampling, the differences between the conditional and unconditional output features progressively shift from being dominated by low-frequency features to high-frequency features. As shown in Fig 14, this observation aligns with the feature visualization analysis: during the early and middle sampling stages, CFG primarily guides the model to synthesize perceptual features such as reasonable shapes and layouts, which are often represented in the low-frequency feature domain. In contrast, during the later stages of sampling, CFG contributes primarily to the synthesis of high-quality details, typically governed by high-frequency features. This insight motivates us to assign higher weights to features of different frequencies at different stages, allowing to gain more emphasis, thereby preserving the visual quality.
Figure 14: The variation in differences between the conditional and unconditional outputs during the sampling process.
A.3.3 FasterCache under different CFG scales and negative prompts
We compared two different negative prompt settings on Open-Sora: (1) default empty negative prompt and (2) non-empty negative prompt:
“worst quality, normal quality, low quality, low res, blurry, text, watermark, logo, banner, extra digits, cropped, jpeg artifacts, signature, username, error, sketch, duplicate, ugly, monochrome, horror, geometry, mutation, disgusting, bad anatomy, bad proportions, bad quality, deformed, disconnected limbs, out of frame, out of focus, dehydrated, disfigured, extra arms, extra limbs, extra hands, fused fingers, gross proportions, long neck, jpeg, malformed limbs, mutated, mutated hands, mutated limbs, missing arms, missing fingers, picture frame, poorly drawn hands, poorly drawn face, collage, pixel, pixelated, grainy”
We calculated the LPIPS, SSIM, and PSNR between the videos generated by FasterCache and those generated by the original model. As shown in Fig. 15 (a) and (b), the experimental results show that FasterCache performs similarly under both prompt settings. This is consistent with our expectations, as CFG-Cache caches the biases between the conditional and unconditional outputs, which are not significantly affected by changes in the negative prompt setting.
We also experimented with different CFG guidance scales $g$ on Open-Sora. As shown in Fig 15 (b) and (c), regardless of increasing or decreasing the scale, while the adjustment affects the original Open-Sora results, FasterCache consistently maintains a high level of alignment with the original results, particularly in preserving details. Therefore, FasterCache is not affected by changes in the CFG guidance scale and maintains high-quality acceleration.
Figure 15: The performance of FasterCache under different CFG scales with empty and non-empty negative prompt settings.
Figure 16: Different Settings of $\alpha$ in CFG-Cache.
A.4 Additional Qualitative Experiments